Nonconstant Growth Valuation

A company currently pays a dividend of $2 per share (D0 = $2). It is estimated that the company's dividend will grow at a rate of 23% per year for the next 2 years, and then at a constant rate of 8% thereafter. The company's stock has a beta of 0.85, the risk-free rate is 7%, and the market risk premium is 3%. What is your estimate of the stock's current price? Do not round intermediate calculations. Round your answer to the nearest cent.
Answer #1

Required rate=risk free rate+Beta*market risk premium

=7+(0.85*3)=9.55%

D1=(2*1.23)=2.46

D2=(2.46*1.23)=3.0258

Value after year 2=(D2*Growth rate)/(Required rate-Growth rate)

=(3.0258*1.08)/(0.0955-0.08)

=210.829935

Hence current price=Future dividend and value*Present value of discounting factor(rate%,time period)

=2.46/1.0955+3.0258/1.0955^2+210.829935/1.0955^2

=$180.44(Approx).